OpenStudy (anonymous):

find the following integral e^(Sqrt(y-2) / sqrt(y-2)

zepdrix (zepdrix):

If we write it like this: \[\Large \int\limits \frac{e^{\sqrt{y-2}}}{\sqrt{y-2}}\;dy\qquad=\qquad \int\limits e^{\sqrt{y-2}}\left(\frac{1}{\sqrt{y-2}}\;dy\right)\]

OpenStudy (anonymous):

oh 2e^(sqrt(y-2)

zepdrix (zepdrix):

Mmm yah that looks right! An extra 2 shows up due to something in there. good job!